Use Chebyshev’s theorem to find what percent of the values will fall between 226 and 340 for a data set with a mean of 283 and standard deviation of 19.
Answer by ewatrrr(10682)

 
Hi,
Use Chebyshev’s theorem to find what percent of the values will fall between 226 and 340 for a data set with a mean of 283 and standard deviation of 19.
226 is (226-283)/19 = 57/19 = 3
340 is (340-283)/19 = 57/19 = 3
According to Chebyshev's Theorem:
"at least [1-(1/3)^2] = [1- 1/9] = 8/9 of the data lies in that interval.
8/9 = 88.89%